OpenStudy (anonymous):

What is the exact distance from (–4, –2) to (4, 6)?

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

I tried to do it by: 8^2 + 4^2 = c^2

OpenStudy (alexandervonhumboldt2):

distance formula is derived from pythogorian theorem

OpenStudy (anonymous):

64+16 = c^2 80 = c^2

OpenStudy (anonymous):

I haven't learned that yet

OpenStudy (anonymous):

answers are 24, 88 100, or 128 all sq roots

pooja195 (pooja195):

\[\huge d=\sqrt{(4--4)^2+(6--2)^2}\] \[\huge d=\sqrt{(4+4)^2+(6+2)^2}\] \[\huge d=\sqrt{(8)^2+(8)^2}\] \[\huge d=\sqrt{64+64}\] \[\huge d=\sqrt{128}\] \[\huge~\rm~d=\sqrt{128}~or~11.31\]

OpenStudy (anonymous):

thanks. how do I give medals? This is my first question.

pooja195 (pooja195):

You may hit the best response button = ) ONLY if you are satisfied with the help.

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

jimthompson5910 (jim_thompson5910):

You can also simplify \(\Large \sqrt{128}\) like so \[\Large \sqrt{128} = \sqrt{64*2}\] \[\Large \sqrt{128} = \sqrt{64}*\sqrt{2}\] \[\Large \sqrt{128} = 8\sqrt{2}\] Notice how I factored 128 into 64*2. One of the factors (64) is a perfect square.